# SynthVLM: High-Efficiency and High-Quality Synthetic Data for Vision Language Models

### v3

@@ 3.2. Synthetic Data Selection
As shown in Figure 3(b), we initially computed CLIPScores for the 1, [^0] K synthetic image-caption pairs. We then selected the top 100K pairs that demonstrated the highest scores, indicating the most accurate and meaningful matches between images and captions. By curating this subset, we constructed a high-quality, highly aligned synthetic dataset.
Table 2. We compared the average CLIPScores of our synthetic dataset, ShareGPT4V, COCO-Caption, and BLIP-LCS. The results indicate that SynthVLM exhibits the highest alignment in terms of CLIPScore values.
Name Sample Avg CLIPScore
COCO-Caption 118K 0.31
BLIP-LCS 558K 0.32
ShareGPT4V 100K 0.32
SynthVLM 1000K 0.34
Curated-SynthVLM 100K 0.38
3.3. High Quality Synthetic Dataset
In this section, we compare commonly used image-caption datasets with the SynthVLM synthetic dataset. The synthetic data offers high image quality, excellent image-text alignment, superior machine ratings, and robust data privacy protection.
High Image Quality. As illustrated in Figure 1, SynthVLM significantly advances image quality by providing images at a resolution of 1024x1024 pixels. This high resolution addresses the common issue of inadequate image quality in existing datasets, thereby supplying high-quality image-caption pairs invaluable for VLLMs. Furthermore, Curated-SynthVLM effectively mitigates issues such as watermarks and advertisements. For image-caption alignment, Curated-SynthVLM leverages the SoTA SDXL to generate images, ensuring that generated images closely correspond to the provided captions.
Excellent Image-Text Alignment. As shown in Table 2, the generated SynthVLM dataset exhibits a higher CLIPScore. By selecting curated image-text pairs of higher quality, Curated-SynthVLM achieves an even higher CLIPScore, surpassing COCO-Caption, BLIP-LCS, and ShareGPT4V. This demonstrates the excellent alignment of our data.
@@ 3.3. High Quality Synthetic Dataset
Sample Model Image-gen win Image-raw win
@@
1K InternVL 692 308
Excellent Machine Rating. Since our data will be used for VLLMs training, we use VLLMs to evaluate the data quality. We selected 1K image-caption pairs and submitted the caption along with the synthetic image and the original image. We used GPT-4 Vision and Intern-VL as the judge model and requested it to select the pair that exhibited higher alignment. The specific prompt used for this evaluation is illustrated in Figure 4. The results, presented in Table 3, demonstrate that images generated have better alignment with the caption.
Protect Data Privacy. A notable advantage of our dataset is its exclusive reliance on captions, which mitigates data privacy concerns. By not using existing images, we ensure that no sensitive or private information associated with those images is compromised. This approach adheres to data privacy best practices, ensuring that our dataset maintains both high quality and ethical integrity.
3.4. SynthVLM
In this section, we utilize the synthetic dataset in section 3.3 to pre-train the Vision Language model. We adopt the same widely-used model architecture summarized in (Zhang et al., 2024; Liu et al., 2023b), as depicted in Figure 5.
Pre-training Stage
As illustrated in Figure 5(a), we train the projector during the pre-training stage to achieve alignment between the image and text modalities. The SynthVLM dataset described in Section 3.3 is utilized for this purpose.
SFT Stage
As shown in Figure 5(b), we further train the projector along with the LLM during the SFT stage to enhance visual understanding capabilities. For this stage, we utilize the commonly used LLaVA 665k dataset from (Liu et al., 2023a).
Through these two training stages, we successfully developed the SynthVLM model. SynthVLM is efficient, utilizing only 100k pre-training data while also protecting privacy by leveraging synthetic data. Additionally, SynthVLM provides a new paradigm for effective alignment between modalities in Vision Language Models using synthetic data.
Figure 5. We introduce the pre-training model structure in Figure (a). Using synthetic data, we pre-train the projector to align the image and text modalities. As shown in Figure (b), we subsequently use LLaVA 665k data to fine-tune the projector and the LLM.
4. Experiment
In this section, we first introduce the experimental setups. We then aim to answer the following questions to verify the effectiveness, efficiency, and privacy protection of our proposed SynthVLM: Q1: Can our SynthVLM achieve SoTA performance compared to previous SoTA methods? Q2: Can our SynthVLM have better image-text alignment compared to previous methods? Q3: How efficient is our SynthVLM compared to previous methods? Q4: Can SynthVLM protect privacy while achieving SoTA performance? Q5: Do we need the generate module and the data quality selection module to enhance model performance?
4.1. Experimental Settings
Datasets.
We utilized the 558k pre-training dataset and the 665k SFT dataset from LLaVA, in addition to the synthetic 100k dataset in section 3.3 for training our SynthVLM.
Models.
For the image generation model, we select SDXL, and SDXL’s steps are set to 60. We employed the well-known LLaVA 1.5 model configured with 13 billion parameters for its robust visual understanding capabilities. For the encoder, we chose the CLIP 336 with a 14-patch configuration, and for the language model, the Vicuna v1.5 configured with 13 billion parameters was selected.
Baselines.
The LLaVA 1.5111https://github.com/haotian-liu/LLaVA model was reproduced using configurations sourced from the official repository to establish a baseline. We opted for the LLaVA model trained with Vicuna v1.5 7B and 13B models as the baseline (Liu et al., 2023b).
Benchmarks.
We select benchmarks for both visual understanding and language understanding. For visual understanding, we choose SQA_Img, MMVet, VizWiz, VQAv2, GQA, MME, and PoPE for a comprehensive evaluation. For pure text benchmarks, we select MMLU and SQA to assess language understanding abilities.
Models LLM SQA SQA_Img MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
Baseline Vicuna-1.5-7B 69.3 67.3 30.5 49.9 78.7 62.5 65.3 1484.8 315.6 86.0 36.3
Synth Select 100k Vicuna-1.5-7B 70.4 68.9 32.2 49.3 79.4 63.1 66.8 1518.5 345.7 87.0 41.2
Baseline Vicuna-1.5-13B 74.2 71.0 35.0 53.6 80.0 63.0 67.7 1531.3 294.5 86.9 52.4
Synth Select 100k Vicuna-1.5-13B 74.9 72.5 35.0 55.9 80.0 63.5 68.3 1573.0 316.1 88.4 54.6
Models LLM SQA MMLU
Avg STEM Humanities Social Sciences Other
Baseline Vicuna-1.5-7B 69.3 36.3 28.6 33.4 39.5 44.5
Synth Select 100k Vicuna-1.5-7B 70.4 41.2 31.7 37.4 47.0 50.2
Baseline Vicuna-1.5-13B 74.2 52.4 41.9 45.8 62.9 61.8
Synth Select 100k Vicuna-1.5-13B 74.9 54.6 45.0 49.3 64.0 62.2
4.2. Synthetic Data Achieves SoTA Performance
To address Q1, we selected LLaVA 1.5 with the CLIP 336 encoder and Vicuna v1.5 7B and 13B as the base model. We used the synthetic 100k dataset to pretrain the LLaVA 1.5 model and subsequently fine-tuned it with LLaVA 665k SFT data. The trained model is denoted as ”Synth Select 100k.” We then compared it with the LLaVA 1.5 model pre-trained on 558k data and fine-tuned on the same 665k SFT data from the official repository, referred to as ”Baseline”
From Table 4, it is evident that Synth Select 100k outperforms the Baseline across all evaluation benchmarks on both the 7B and 13B model. Specifically, Synth Select 100k achieves SoTA results on visual benchmarks such as SQA_Img, MMVet, VizWiz, VQAv2, GQA, MME, MMB, and PoPE. Furthermore, SynthVLM also excels in pure language benchmarks, demonstrating superior performance in SQA and MMLU, thus showcasing its comprehensive capabilities in both vision and language tasks.
4.3. Effective Vision Language Alignment
To address Q2, we utilize pure language ability to demonstrate the model’s alignment. During the SFT stage, the LLM can be adjusted to align with images; hence, better alignment can preserve the LLM’s pure language ability (Wang et al., 2023). We select the MMLU and SQA benchmarks for their comprehensive language understanding capabilities.
From Table 5, it is evident that our Synth Select 100k model outperforms the Baseline on all MMLU benchmarks and SQA benchmarks.
@@ 4.4. Efficient Vision Language Alignment
To address Q3, we examine the computational resource usage during training and evaluate data utilization efficiency for generating image-caption pairs.
@@
As shown in Table 6, by integrating a data selection module, our approach utilizes only 19% of the LLAVA data and 10% of the original synthetic data while achieving SoTA performance. This demonstrates that our data selection method can reduce computational usage by more than 80%.
@@
Overall, our method efficiently aligns image and text modalities, demonstrating strong potential for effective modality alignment. Additionally, acquiring captions is considerably less resource-intensive than obtaining images. By integrating existing large language models (LLMs), it is feasible to fully automate the generation of both captions and images, further enhancing the efficiency of dataset construction in VLLMs.
4.5. Privacy Protection Pre-training
To address Q4, we compare the synthetic image and the original image in Figure 6 and Figure 7. We can see synthetic data offers significant advantages in protecting data privacy.
As illustrated in Figure 6, synthetic image (a) effectively avoids privacy issues by not representing real human faces, while original image (b) contains human faces, potentially leading to privacy concerns. Similarly, in Figure 7, synthetic images in (a) show vehicles and tickets without revealing real license plates and ticket information, ensuring privacy protection. In contrast, original images in (b) display actual license plates and ticket information, which can potentially lead to privacy issues.
By utilizing generative models like DDPM, synthetic data can be created with similar statistical properties to real data without involving actual personal information. This provides a secure pathway for data sharing, model training, and analysis, helping to comply with privacy regulations, protecting user privacy, and simultaneously advancing the fields of artificial intelligence and machine learning.
@@ 4.5. Privacy Protection Pre-training
Models LLM SQA SQA_Img MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
Synth Select 100k Vicuna-1.5-7B 70.4 68.9 32.2 49.3 79.4 63.1 66.8 1518.5 345.7 87.0 41.2
w/o generation 100k Vicuna-1.5-7B 69.3$\downarrow$ 67.0$\downarrow$ 31.2$\downarrow$ 46.8$\downarrow$ 79.3$\downarrow$ 62.9$\downarrow$ 66.2$\downarrow$ 1488.8$\downarrow$ 327.5$\downarrow$ 86.2$\downarrow$ 39.1$\downarrow$
w/o selection 100k Vicuna-1.5-7B 69.9$\downarrow$ 67.7$\downarrow$ 30.2$\downarrow$ 50.2 79.1$\downarrow$ 62.2$\downarrow$ 63.5$\downarrow$ 1421.7$\downarrow$ 301.8$\downarrow$ 87.3 40.6$\downarrow$
Synth Select 100k Vicuna-1.5-13B 74.9 72.5 35.0 55.9 80.0 63.5 68.3 1573.0 316.1 88.4 54.6
w/o generation 100k Vicuna-1.5-13B 73.6$\downarrow$ 71.4$\downarrow$ 33.0$\downarrow$ 53.6$\downarrow$ 80.0 63.4$\downarrow$ 67.5$\downarrow$ 1514.3$\downarrow$ 295.7$\downarrow$ 88.2$\downarrow$ 53.6$\downarrow$
w/o selection 100k Vicuna-1.5-13B 74.1$\downarrow$ 70.5$\downarrow$ 35.6 53.2$\downarrow$ 79.7$\downarrow$ 63.1$\downarrow$ 67.5$\downarrow$ 1512.7$\downarrow$ 303.2$\downarrow$ 86.9$\downarrow$ 53.0$\downarrow$
@@
Figure 6. From (a), it is evident that synthetic images can avoid representing real human faces, while (b) contains human faces, potentially leading to privacy issues.
@@ 4.6. Ablation Study
To address Q5, we conducted the following ablation study. Specifically, we conducted an ablation study where we removed the data generation module and the data selection module separately to evaluate their individual contributions to the effectiveness of our data generation pipeline.
@@
The exclusion of the data generation module significantly impacts the model’s performance, as illustrated in Tables 7 and 8, labeled as ”w/o generation 100k”. The variant without this module demonstrates markedly lower accuracy across all benchmarks. These results emphasize the crucial role of the data generation process in sustaining the high performance of the SynthVLM model. This also underscores SynthVLM’s potential in constructing highly aligned datasets.
@@
The absence of the data selection module similarly leads to a noticeable decline in performance, indicated as ”w/o selection 100k” in Tables 7 and 8. Given the inherent randomness of diffusion models, which inevitably generate some low-quality images, the data selection module is crucial for removing these subpar elements.
Overall, the ablation study highlights the critical role of data generation and data selection modules in SynthVLM. These experiments provide valuable insights into the contributions of each module, guiding future improvements and optimizations of the SynthVLM model.
@@ 5. Conclusion
In recent years, with the development of VLLMs, data generation has become increasingly important. Synthetic images are crucial for LLMs due to the lack of high-quality data for extensive training. In this paper, we propose a new image generation pipeline for generating VLLMs pre-training data, providing a new paradigm for image generation for VLLMs. Remarkably, the synthetic data trained SynthVLM model outperforms the baseline using only 18% synthetic data. Additionally, it achieves SoTA alignment ability efficiently. Furthermore, SynthVLM protects data privacy by using synthetic datasets.

### v4

@@ 3.2. Step2: Synthetic Data Selection
As shown in Figure 1(b), we initially computed CLIPScores and SSIMScore for the 1M synthetic image-caption pairs. We then selected the top 100K pairs that demonstrated the highest scores, indicating the most accurate and meaningful matches between images and captions. By curating this subset, we constructed a high-quality, highly aligned synthetic dataset.
@@
Sample Model Sythetic Image win Web Image win
@@
1K InternVL2 692 308
1K Human 758 242
Table 3. We compared the average CLIPScore, SSIMScore, and their weighted score across our synthetic dataset, ShareGPT4V, COCO-Caption, and BLIP-LCS. The results demonstrate that SynthVLM achieves the highest alignment and image quality.
Datasets # Samples CLIPScore SSIMScore Weighted_Score
COCO-Caption 118K 0.31 0.73 0.67
BLIP-LCS 558K 0.32 0.75 0.70
ShareGPT4V 100K 0.32 0.79 0.71
Synth Dataset 1000K 0.34 0.78 0.73
Curated-Synth Dataset 100K 0.36 0.86 0.79
@@
Models LLM SQA SQAI MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
LLaVA-7B Vicuna-1.5-7B 69.3 67.3 30.5 49.9 78.7 62.5 65.3 1484.8 315.6 86.0 36.3
SynthVLM-7B Vicuna-1.5-7B 70.4 68.9 32.2 49.3 79.4 63.1 66.8 1518.5 345.7 87.0 41.2
LLaVA-13B Vicuna-1.5-13B 74.2 71.0 35.0 53.6 80.0 63.0 67.7 1531.3 294.5 86.9 52.4
SynthVLM-13B Vicuna-1.5-13B 74.9 72.5 35.0 55.9 80.0 63.5 68.3 1573.0 316.1 88.4 54.6
Models LLM SQA MMLU
Avg STEM Humanities Social Sciences Other
LLaVA-7B Vicuna-1.5-7B 69.3 36.3 28.6 33.4 39.5 44.5
SynthVLM-7B Vicuna-1.5-7B 70.4 41.2 31.7 37.4 47.0 50.2
LLaVA-13B Vicuna-1.5-13B 74.2 52.4 41.9 45.8 62.9 61.8
SynthVLM-13B Vicuna-1.5-13B 74.9 54.6 45.0 49.3 64.0 62.2
3.3. High Quality Synthetic Dataset
In this section, we compare commonly used image-caption datasets with the SynthVLM-100K dataset. The synthetic data offers high image quality, excellent image-text alignment, superior model ratings, and robust data privacy protection.
High Image Quality
As illustrated in Figure 2, SynthVLM markedly improves image quality by generating images at a resolution of 1024x1024 pixels. This high resolution addresses the prevalent issue of suboptimal image quality in existing datasets, providing high-quality image-caption pairs that are invaluable for training VLLMs. Additionally, SynthVLM-100K effectively eliminates artifacts such as watermarks and advertisements.
Excellent Image-Text Alignment.
As shown in Table 3, the SynthVLM-100K dataset exhibits a higher CLIPScore and SSIMScore compared to existing high-quality web datasets. By selecting curated image-text pairs of superior quality, SynthVLM-100k surpasses datasets such as COCO-Caption, BLIP-LCS, and ShareGPT4V. This demonstrates the outstanding alignment of our dataset.
Excellent Model Rating.
Since our data will be used for VLLMs training, we use VLLMs to evaluate the data quality. We selected 1K image-caption pairs and submitted the caption along with the synthetic image and the original image. We used human annotators, GPT-4 Vision and Intern-VL2 as the judge model and requested it to select the pair that exhibited higher alignment. The specific prompt used for this evaluation is illustrated in Figure 3. The results, presented in Table 2, demonstrate that images generated have better alignment with the caption.
Table 7. The results on the MME benchmark demonstrate that using generated data can still maintain a leading performance in real-world problems, further expanding the application scope of SynthVLM.
Models LLM MME Cognition MME Perception
Reasoning Translation Code Posters Celebrity Scene Landmark Artwork OCR
LLaVA1.5-7B Vicuna-1.5-7B 126.4 57.5 62.5 148.3 132.1 143.0 141.8 123.8 100.0
SynthVLM-7B Vicuna-1.5-7B 143.6 70.0 57.5 157.1 124.2 157.8 156.5 136.2 102.5
LLaVA-13B Vicuna-1.5-13B 119.3 50.0 62.5 155.4 127.4 158.5 165.2 129.3 110.0
SynthVLM-13B Vicuna-1.5-13B 130.0 65.0 62.5 160.4 138.2 162.0 158.8 140.3 104.5
Protect Data Privacy.
A notable advantage of our dataset is its exclusive reliance on captions, which mitigates data privacy concerns. By not using existing images, we ensure that no sensitive or private information associated with those images is compromised. This approach adheres to data privacy best practices, ensuring that our dataset maintains both high quality and ethical integrity.
Through these two training stages, we successfully developed the VLLM model trained by SynthVLM-100K data. Our method is efficient, utilizing only 100K pre-training data while also protecting privacy by leveraging synthetic data. Additionally, SynthVLM provides a new paradigm for effective alignment between modalities in Vision Language Models using pure synthetic data.
4. Experiments
In this section, we utilize the image data synthesis system for various tasks. We then aim to answer the following questions to verify the effectiveness, efficiency, and privacy protection of our proposed SynthVLM: Q1: Can our SynthVLM-7B and SynthVLM-13B achieve SOTA performance compared to previous SOTA methods? Q2: How does the efficiency of our SynthVLM compare to previous methods? Q3: Can SynthVLM protect privacy while achieving state-of-the-art performance? Q4: Would using generated data in full impact the application of the model in real-world scenarios? Q5: Do we need the generate module and the quality selection module to enhance model performance?
4.1. Experimental Settings
Models.
We use the LLaVA 1.5 model to validate the effectiveness of our dataset. For the Vision Encoder, we select CLIP 336px, and for the LLM, we use Vicuna 7B and Vicuna 13B. Our training parameters are consistent with LLaVA, and we adopt a two-stage training approach. In the pre-training stage, we train the projector to align the image and text modalities. The SynthVLM-100K dataset described in Section 3.3 is used for this purpose. In the SFT stage, we further train the projector alongside the LLM to improve visual understanding capabilities.
Datasets.
For LLaVA-BaseLine-7B and LLaVA-BaseLine-13B, we use the LLaVA 558k dataset for pre-training and the LLaVA 665k dataset for SFT. For SynthVLM-7B and SynthVLM-13B, we use the SynthVLM-100K dataset for pre-training and the LLaVA 665k dataset for SFT.
Benchmarks.
We select benchmarks for both visual understanding and language understanding. For visual understanding, we choose SQAI, MMVet, VizWiz, VQAv2, GQA, MME, and PoPE for a comprehensive evaluation. For language benchmarks, we select MMLU and SQA to assess language understanding abilities.
Settings.
We primarily adhered to the hyperparameters specified in the official repository for the LLaVA model evaluations. All experiments were conducted on an 8x NVIDIA A100 GPU machine equipped with a 120-core CPU and 960GB of memory.
4.2. Synthetic Data Achieves SOTA Performance
To address Q1, we trained the model described in experimental settings. From Table 5, it is evident that SynthVLM model outperforms the Baseline across all evaluation benchmarks on both 7B and 13B. Specifically, SynthVLM model achieves SOTA results on visual benchmarks such as SQAI, MMVet, VizWiz, VQAv2, GQA, MME, MMB, and PoPE. Furthermore, SynthVLM model also excels in pure language benchmarks, demonstrating superior performance in SQA and MMLU, thus showcasing its comprehensive capabilities in both vision and language tasks.
@@ 4.3. Efficient Vision Language Alignment
To address Q2, we examine the computational resource usage during training and evaluate the data utilization efficiency for generating image-caption pairs.
@@
By integrating a data selection module, our approach utilizes only 19% of the LLAVA data and 10% of the original synthetic data while achieving SOTA performance. This demonstrates that our data selection method can reduce computational usage by more than 80%.
@@
Models LLM SQA SQAI MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
SynthVLM-7B Vicuna-1.5-7B 70.4 68.9 32.2 49.3 79.4 63.1 66.8 1518.5 345.7 87.0 41.2
w/o generation Vicuna-1.5-7B 69.3$\downarrow$ 67.0$\downarrow$ 31.2$\downarrow$ 46.8$\downarrow$ 79.3$\downarrow$ 62.9$\downarrow$ 66.2$\downarrow$ 1488.8$\downarrow$ 327.5$\downarrow$ 86.2$\downarrow$ 39.1$\downarrow$
w/o selection Vicuna-1.5-7B 69.9$\downarrow$ 67.7$\downarrow$ 30.2$\downarrow$ 50.2 79.1$\downarrow$ 62.2$\downarrow$ 63.5$\downarrow$ 1421.7$\downarrow$ 301.8$\downarrow$ 87.3 40.6$\downarrow$
SynthVLM-13B Vicuna-1.5-13B 74.9 72.5 35.0 55.9 80.0 63.5 68.3 1573.0 316.1 88.4 54.6
w/o generation Vicuna-1.5-13B 73.6$\downarrow$ 71.4$\downarrow$ 33.0$\downarrow$ 53.6$\downarrow$ 80.0 63.4$\downarrow$ 67.5$\downarrow$ 1514.3$\downarrow$ 295.7$\downarrow$ 88.2$\downarrow$ 53.6$\downarrow$
w/o selection Vicuna-1.5-13B 74.1$\downarrow$ 70.5$\downarrow$ 35.6 53.2$\downarrow$ 79.7$\downarrow$ 63.1$\downarrow$ 67.5$\downarrow$ 1512.7$\downarrow$ 303.2$\downarrow$ 86.9$\downarrow$ 53.0$\downarrow$
@@
4.4. Privacy Protection Synthetic Data
To address Q3, we compare the synthetic image and the original image in Figure 4. Synthetic data offers significant advantages in protecting data privacy. In Figure 4, synthetic images in (a) show vehicles and tickets without revealing real license plates and ticket information, ensuring privacy protection. In contrast, original images in (b) display actual license plates and ticket information, which can potentially lead to privacy issues.
4.5. Capability to Solve Real-World Problems
A critical issue when using generated data is whether the model loses its ability to solve real-world problems. To address Q4, we utilized the MME Benchmark to assess the model’s performance in various real-world scenarios, such as artwork, celebrity, code reasoning, landmarks, and posters. The specific results are presented in Table 8. Using SynthVLM-100K, our model performs consistently well, matching or even surpassing the baseline across most real-world tasks. We attribute this success to the diversity of the captions and the generalization ability of the Diffusion model.
@@ 4.6. Ablation Study
To address Q5, we conducted the following ablation study. Specifically, we conducted an ablation study where we removed the data generation module and the data selection module separately to evaluate their individual contributions to the effectiveness of our data generation pipeline. In this section, we controlled the number of samples in each experimental group to be 100K.
@@
The exclusion of the data generation module significantly impacts the model’s performance, as illustrated in Tables 8 and 9, labeled as ”w/o generation”. The variant without this module demonstrates markedly lower accuracy across all benchmarks. These results emphasize the crucial role of the data generation process in sustaining the high performance of the SynthVLM model. This also underscores SynthVLM’s potential in constructing highly aligned datasets.
@@
The absence of the data selection module similarly leads to a noticeable decline in performance, indicated as ”w/o selection” in Tables 8 and 9. Given the inherent randomness of diffusion models, which inevitably generate some low-quality images, the data selection module is crucial for removing these subpar elements.
Overall, the ablation study highlights the critical role of the data generation and the data selection modules in SynthVLM. These experiments provide valuable insights into the contributions of each module, guiding future improvements and optimizations of the SynthVLM model.
@@ 5. Conclusion
Data is crucial for training VLLMs. However, high-quality web data is limited. In this paper, we introduce a novel image synthesis pipeline to generate pre-training data for VLLMs. Our synthetic images are of higher quality as they are free from watermarks and advertisements. Furthermore, our synthetic image-caption pairs exhibit superior CLIPScore and SSIMScore, reflecting better image-caption alignment and image quality. Remarkably, the SynthVLM model, trained with only 18% synthetic data, surpasses the LLaVA baseline model trained on the full dataset. Moreover, it efficiently achieves SOTA alignment capabilities while preserving the language proficiency of VLLMs. By employing synthetic datasets, SynthVLM also safeguards data privacy, underscoring the practicality and effectiveness of our data synthesis pipeline.